OpenStudy (anonymous):

help me;-) what is 26^6 X3 to the fifth power i need help ;)

OpenStudy (anonymous):

helppp

OpenStudy (mathstudent55):

When you raise a power to a power, you multiply the powers. \(\large (26^6 \times 3)^5 = 26^{6 \times 5} \times 3^5 = 26^{30} \times 3^5\)